Subject: Minutes of Sep. 8 APO 3.5m User's Committee Meeting

From: strauss@astro.Princeton.EDU

Submitted: Wed, 9 Sep 1998 12:12:15 -0400 (EDT)

Message number: 305 (previous: 304, next: 306 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************September 8, 1998**********************

Attending: Michael Strauss, Bruce Gillespie, Ed Turner, Jeff Brown,
Chris Stubbs, Alan Uomoto, Jon Holtzman. 

Missing: Ed Kibblewhite

  The shutdown tasks were mostly for repairs that make things
more reliable and easier to maintain; most of these were successfully
accomplished.  There was a plan to clean and resurface the drive boxes,
especially the altitude drive, but this was not done due to lack of
time and personnel.  This will be done during scheduled engineering
time in the fourth quarter.  Interestingly, during the shutdown, the
weather was so bad anyway, that only roughly 3 clear nights were lost
in July.  

  Status of GRIM2: 
  It shows some flaky behavior; every once in a
while, it returns a screwball frame, as if there are electronic
glitches.  There is a fear that the GRIM could fail catastrophically
at any time, and could not be brought back to life.  It is also no
longer a state of the art instrument, and is no longer competitive
with IR imaging instruments at other sites.  Finally, the IR expertise
that built GRIM2 has mostly left the ARC consortium, and their
expertise has not been transferred to the APO staff, so it is
difficult to maintain the instrument.  There have been stated plans
for upgrading the instrument (for example, so that it can do multiple
reads), but these have never gotten off the ground.  In the meantime,
the bright time is being more and more under-subscribed.  Turner and
Gillespie are looking into the possibility of getting new IR
instruments from outside the consortium, as there is not any obvious
expertise/interest within the consortium for building such an
instrument (or just keeping GRIM2 alive).

Observing Specialists: 
  Camron Hastings has accepted an offer to become a full-time
observing specialist.  We're up to our full complement of 3. 

  As an experiment, we are trying a new scheduling of observing
specialists: most nights, the night will be split at midnight between
2 observing specialists.  This allows more continuity between
afternoon setups and observing, and means that observing specialists
won't have impossibly long runs in the winter.

Rotator problems: 
  The rotator has been mis-behaving quite a bit over the last week.
During shutdown, lots of guider mechanicals were overhauled.  As a
consequence, the rotator movement is much more smooth.  The mechanical
stiction is possibly not longer well-matched to the servo motors that drive it.
Also, the drive amplifier that drives the rotator was rebuilt; there
seems to be some real problems with this new, "improved" amplifier.
This has been swapped out now, and sent back to the manufacturer;
there is hope that the rotator should behave better now.  The APO
staff continue to work on it. 

Status of Secondary: 
  As of last week, f/9 secondary for MMT was almost complete at
Steward; we are next!  Thus our place in the queue has not slipped.

  News from Echelle: There will be a meeting at Chicago, September 17,
between Turner, Gillespie, Stubbs, Jeff Morgan, Ed Jenkins, and the
people building the echelle.  They will discuss the current status of
the instrument and its software, and the schedule for bringing the
instrument to the mountain.

  SPICAM had an electronics failure in the dewar recently.  It was
sent back to UW, where it was repaired, the cryogenics is being
revamped, replacing the LN2 system with a mechanical "cryo-tiger"
system.  The goal is to get it back in operation for the upcoming dark
run.  In addition, there will be some upgrades to electronics and
software, including faster clears, and perhaps more flexibility in the
read-out time/read noise trade-off.

  DIS proposal: Magnier and Stubbs of UW have put in a proposal to NSF
to do upgrades to DIS: in particular, to install new detectors with
better sampling and wavelength coverage, and improved read
electronics.  Stubbs is fairly optimistic about this proposal. 

  Long-term plans: we need new instruments!  User's committee stands
strongly behind this statement. 


  A lot of tree thinning has been happening in the forests nearby
recently, to improve tree's health (it reduces the spread of
diseases), and to reduce fire danger.  The Forest Service burns the
left-over branches, etc.; creating a real danger of falling ash.
There have been discussions with the Forest Service to do these burns
only once every few months, at previously scheduled times (say, at
full moon, when engineering work can be scheduled), rather than

  The new electronic web schedule is working out nicely.  Check it
out, off the APO home page; and send suggestions to Craig Loomis.
There is a trade form there as well, useful for people to exchange
time with one another. 

  The number of people who have expressed interest in the APO User's
Community meeting in Seattle October 23-24 is quite small (and
consists almost entirely of people from Seattle and Pullman!); the
User's Committee will work to drum up more interest among their
respective departments.  The previous two such meetings have been very
successful, and have helped us decide the directions we want to go
with the observatory. 

  In this respect, we discussed the highest long-term priority, which
is new and state of the art instrumentation for the observatory.  What
this really requires is a mechanism in place to facilitate such
instruments being built by the consortium.  The User's Committee will
draft a brief letter to the director, stating that this is their

  Last month's minutes are approved.

  Next meeting, Monday, September 28, 12:30 PM

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 305 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to